# UniGenX: Unified Generation of Sequence and Structure with Autoregressive Diffusion

### v1

@@ 5.2 Molecular Results
As shown in Table 3 and Table 4, the conformations generated by our model are able to achieve good Coverage and Match score whether based on recall or precision. Overall, the results are comparable to DMCG. It is noticeable that for every metric’s median we outperform the baselines; the remaining ones are only a little lower than DMCG [36].
@@
Figure 5: Comparison of energy (A, C) and gap (B, D) property distributions between ground truth and generated conformations. A and C, and B and D, represent alternative visualizations of the same distributions.
The results of property prediction are shown in Table 5. Our method significantly outperforms RDKit, GeoMol, ConfGF and DMCG, which shows the effectiveness of our method. Comparison between the distribution of energies and gaps of ground truth and generated conformations is in Figure 5.
@@
Units Bohr3 meV meV meV D $\frac{\text{cal}}{\text{mol}}$
@@
Table 6 presents the LDM sampling results for conditional generation as introduced in [62], where UniGenX achieves state-of-the-art (SOTA) performance across all five properties. However, we believe there is an issue with the evaluation results provided by [62]. Following [61], when we used another portion of the QM9 dataset to assess the performance of the classifier from [62], we found that the values for $\epsilon_{H}$, $\epsilon_{L}$ and $\Delta\epsilon$ were approximately 27 times larger than those reported in [61]. Given that $1Ha=27.2114eV$, we suspect there is a unit discrepancy in [62]. While the unit in their paper is listed as meV, we believe the correct unit should be mHa. We highlight this error here and have corrected it in our table.
@@
Table 7 presents the conditional generation results using EDM sampling, as introduced in [61]. UniGenX achieves state-of-the-art (SOTA) performance on four out of six properties. Notably, for the property $\mu$, our model demonstrates a significant improvement of $53.6\%$ over the previous SOTA [70]. However, performance on $\alpha$ and $C_{v}$ is slightly lower compared to [69]. This can be attributed to two factors: first, [69] employs separate models for each property, whereas UniGenX uses a unified model, potentially leading to performance trade-offs. Second, [69] incorporates post-generation molecular force field optimization, following [62], which leverages domain-specific expert knowledge. Our approach evaluates the model’s direct output without such optimization. While these differences may impact direct comparison, UniGenX still excels in key performance metrics, highlighting its versatility and robustness.
@@
The properties in Table 7 were calculated for generated samples using an EGNN model, consistent with the EDM settings from [61]. The QM9 row indicates the EGNN model’s error estimation relative to the QM9 property labels. Notably, UniGenX’s generated results exhibit error closely approximating the validation model’s error. For instance, the HOMO gap is identical to QM9 (39 vs. 39), and the LUMO gap is lower (35 vs. 36). This suggests that UniGenX’s performance may exceed the validation model’s capacity for accurate error assessment. A detailed discussion of this discrepancy is reserved for future work.
@@
Figure 6: Examples of groundtruth (left) and generated (right) molecular structures
@@
Methods MR(%$\uparrow$ RMSD $\downarrow$ MR(%$\uparrow$ RMSD $\downarrow$ MR(%$\uparrow$ RMSD $\downarrow$
@@ 5.3 Intergrating with LLM Results
Methods MR(%$\uparrow$ RMSD $\downarrow$
UniGenX-NL(1B) 63.10 0.0397
Table 10: Results of UniGenX with capability of natural langauge.
The results shown in 10 demonstrate that our model retains its linguistic capabilities without compromising accuracy, providing a solid foundation for future work, such as integrating language with a unified model for both linguistic and numerical tasks.
@@ 6 Conclusion
Appendix A Diffusion Model and Flow Matching
Theorem 1.
$\min_{Z_{\theta}\in\mathcal{V}}\|Z_{\theta}-X\|\,,$
[…]
A.1 Relationship Between Diffusion Models and Flow Matching
$x_{t}=\alpha_{t}x+\sigma_{t}z,$ (16)
$\frac{dx_{t}}{dt}=\dot{\alpha}_{t}x+\dot{\sigma}_{t}z,$ (17)
A.2 Connections of the Training Targets
@@ Appendix B Pseudocode
2:$w_{gt}$: Ids of input tokens in the input sentences
3:$v_{gt}$: Numeral values in the input sentences
4:$m_{val}$: Masks for values
5:$m_{pad}$: Padding Masks
@@
10:$x_{input}[\neg m_{v}]\leftarrow$ Embedding($w$)
11:$x_{input}[m_{v}]\leftarrow$ Linear($v$)
@@
13:$h_{rep}\leftarrow$ Transformer_Decoder($x_{input},m_{pad}$)
14:$w_{logits}\leftarrow$ Linear($h_{rep}[\neg m_{val}]$)
@@
2:$v_{gt}$: Ground truth numeral values
3:$h_{rep}$: Hidden representations for conditional diffusion
@@
7:$x_{t}\leftarrow\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-{\overline{\alpha}_{t}}}\epsilon$
@@
2:$w_{in}$: Ids of input tokens in the initial sentences
3:$v_{in}$: Numeral values in the initial sentences
4:$m_{val}$: Masks for values
5:$m_{pad}$: Padding Masks
6:$\texttt{Initialize}~{}w\leftarrow w_{in},v\leftarrow v_{in}$
@@
[…]
@@
8:$r\leftarrow\texttt{diff}.\texttt{p\_sample\_loop}(\texttt{net},y,model\_kwargs)\textbf{ \# Predicted denoised sample}$
@@ C.1 Dictionary for Diverse Domains and Tasks
The vocabulary dictionary also incorporates special tokens, which are crucial for UniGenX’s flexibility as a next-token prediction model across diverse domains. These tokens serve as domain indicators, such as $<$bof$>$ for material formulas and $<$bos$>$ for molecule SMILES. Extending UniGenX to other domains, like proteins and DNA, requires adding corresponding tokens, such as $<$boa$>$ and $<$bod$>$. This design empowers UniGenX to handle various domains within a unified sequence representation. Thus, these domain-specific special tokens are integral to the vocabulary dictionary.
Furthermore, special tokens facilitate different tasks. For structure generation, $<$boc$>$ denotes coordinates. For conditional generation of sequences and structures, property tokens like $<$bulk$>$ for bulk modulus and $<$band$>$ for band gap are used. These tokens, followed by property values, precede the molecule sequence and coordinates. This allows UniGenX to generate both sequences and structures conditioned on specific properties. Therefore, the special tokens in the vocabulary dictionary provide UniGenX with flexibility across domains and tasks.
@@ C.5 Conditional Generation of Molecules
C.6 Instruction Tuning with Natural Language Model

### v2

@@ 4.2 Molecular Results
As shown in Table 4 and Table 5, the conformations generated by our model are able to achieve good Coverage and Match score whether based on recall or precision. Overall, the results are comparable to DMCG [20]. It is noticeable that for every metric’s median we outperform the baselines; the remaining ones are only a little lower than DMCG.
@@
Figure 9: Comparison of energy (A) and gap (B) property distributions between ground truth and generated conformations.
The results of property evaluation are shown in Table 6. Our method significantly outperforms SOTA and baseline methods listed in the Table, which shows the effectiveness of our method. Comparison between the distribution of energies and gaps of ground truth and generated conformations is in Figure 9.
@@
Units Bohr3 meV meV meV D $\frac{\text{cal}}{\text{mol}}$K
@@
Table 7 presents sampling results according to equation 5 for conditional generation as introduced in [28], where UniGenX achieves state-of-the-art (SOTA) performance across all five properties. Compared with the latest state-of-the-art methods, UniGenX achieves up to a 260% improvement and at least a 34% improvement across various evaluation metrics.
Remark 2.
@@
Corrected the typo from meV to Hain in Table 7. There is a typo with the evaluation results provided by [28]. Following [27], when we used another portion of the QM9 dataset to assess the performance of the classifier from [28], we found that the values for $\epsilon_{H}$, $\epsilon_{L}$ and $\Delta\epsilon$ were approximately 27.21 times larger than those reported in [27]. Given that $1Ha=27.2114eV$, we suspect there is a unit discrepancy in [28]. While the unit in their paper is listed as meV, we believe the correct unit should be mHa. Confirmed with the authors in the papers and we corrected it in our Table 7.
@@
Table 8 presents the results sampling as given in equation 6 , as introduced in [27]. UniGenX achieves state-of-the-art (SOTA) performance on four out of six properties. Notably, for the property $\mu$, our model demonstrates a significant improvement of $53.6\%$ over the previous SOTA [46]. However, performance on $\alpha$ and $C_{v}$ is slightly lower compared to [45]. This can be attributed to two factors: first, [45] employs separate models for each property, whereas UniGenX uses a unified model, potentially leading to performance trade-offs. Second, [45] incorporates post-generation molecular force field optimization, following [28], which leverages domain-specific expert knowledge. Our approach evaluates the model’s direct output without such optimization. While these differences may impact direct comparison, UniGenX still excels in key performance metrics, highlighting its versatility and robustness.
@@
The properties in Table 8 were calculated for generated samples using an EGNN model, consistent with the EDM settings from [27]. The QM9 row indicates the EGNN model’s error estimation relative to the QM9 property labels. Notably, UniGenX’s generated results exhibit error closely approximating the validation model’s error. For instance, the MAE of HOMO energy level is identical to QM9 (39meV vs. 39meV), and the MAE of LUMO is lower (35meV vs. 36meV). This suggests that UniGenX’s performance may exceed the validation model’s capacity for accurate error assessment. A detailed discussion of this discrepancy is reserved for future work.
For the multiple-property generation task, we visualize the results in the following way: for a single property, we plot a kernel density histogram 10(a); for two properties, we use a scatter plot 10(b); and for three or more properties, we apply principal component analysis (PCA) to reduce the dimensions and plot a 2D scatter plot 11. These visualizations demonstrate that our model can flexibly generate molecules conditioned on varying numbers of properties, while maintaining stability and reliability.
(a) Kernel density estimate (KDE) plot comparing the distribution of the given target property and the predicted property values of generated molecules.
(b) Scatter plot showing the joint distribution of two predicted properties for generated molecules conditioned on given target properties.
Figure 11: PCA projection of predicted properties of generated molecules conditioned on three or more target properties.
@@
Figure 12: Examples of predicted structures. The groundtruth and predicted structures are shown side-by-side, with groundtruth on the left, generated structures on the right.
@@
4.3 Protein Results
Figure 13: Free energy surfaces of DESRES fast-folding proteins. The landscapes on the left are obtained from long-timescale molecular dynamics (MD) simulations, while the landscapes on the right are generated by UniGenX. The model successfully captures the key features of the energy landscapes, including native and unfolded states, with an average deviation of only 0.91 kcal/mol from the MD reference—a discrepancy comparable to that between different classical force fields [47, 48]. Representative conformational states predicted by UniGenX are also shown.
As shown in Figure 13, for all proteins, our model successfully predicts native and unfolded states with similar topological features on the free energy landscape. The predicted two-dimensional free energy surfaces closely resemble those obtained from MD simulations. Notably, our model captures partial or complete folding intermediates that are visible on the landscape, demonstrating its capability to approximate the equilibrium distribution. Quantitatively, the average discrepancy between our model and molecular dynamics (MD) simulations on the 2D free energy landscapes is 0.91 kcal/mol, ranging from 0.64 kcal/mol for 1FME to 1.20 kcal/mol for A3D. Notably, this level of deviation is comparable to the expected differences between two classical MD force fields [47, 48], suggesting that UniGenX achieves near-physically consistent accuracy.
As shown in Table 9, UniGenX is able to generate protein sequences that not only match the target EC number but also exhibit strong structural quality and functional relevance. When comparing UniGenX with and without structural conditioning during training, we observe consistent improvements across both sequence-level and structure-level metrics. Specifically, models trained with structural information yield higher pLDDT and pTM scores, along with significantly better TM-scores and lower RMSDs when benchmarked against both AFDB and PDB reference structures.
Moreover, comparison with structures predicted by ESMFold (Table 10) reveals a high degree of structural agreement: generated sequences consistently achieve TM-scores above 0.85 and low RMSD values, indicating that UniGenX reliably produces proteins that fold into plausible and stable 3D structures aligned with the intended enzymatic function.
Figure 14 presents one representative case for each of the three tested EC numbers. Notably, despite relatively low sequence identity to naturally occurring proteins, the generated sequences fold into structures with strong similarity to experimentally resolved counterparts. This demonstrates that UniGenX can generalize beyond sequence similarity, capturing structural patterns that are functionally meaningful and biophysically consistent.
Tested EC number Overview Sequence
pLDDT pTM Time(s) Accuracy (%) Avg. Top Hit Identity (%)
UniGenX 2.6.1.x 80.11 0.821 9.38 63.87 77.35
3.1.1.x 87.25 0.883 3.77 76.42 75.31
3.5.2.x 91.59 0.938 8.28 84.95 77.49
w/o struct. 2.6.1.x 74.96 0.766 9.20 69.48 71.58
3.1.1.x 85.83 0.878 3.37 74.28 70.07
3.5.2.x 87.46 0.911 9.13 79.91 71.84
Table 9: Evaluation of function-guided enzyme generation. The performance of UniGenX trained with both sequence and structure is compared to a sequence-only baseline. Including structural information during training leads to generated proteins with higher predicted structural quality (pLDDT, pTM) and functional accuracy.
Tested EC Number RMSD (Å) TM-score
2.6.1.x 4.398 0.916
3.1.1.x 6.047 0.867
3.5.2.x 3.206 0.936
Table 10: Structural comparison between UniGenX generated structures and ESMFold-predicted structures across three EC numbers.
Figure 14: For all three cases, the proteins generated by UniGenX exhibit the close structural alignment to naturally occurring proteins with relative small sequence identity.
4.4 Protein-Ligand Docking Results
We benchmark our proposed UniGenX (100M) model against a modified version of 3DMolformer[33], in which the diffusion module is replaced with a deterministic linear transformation. To ensure a fair comparison, both models adopt the same Transformer backbone and are trained under identical hyperparameters for 1,000 epochs using a cosine learning rate schedule. Quantitative results are presented in Table 11.
To complement these aggregate metrics with a more nuanced view of performance, Figure 15 shows the cumulative distribution functions (CDFs) of RMSD values for predicted ligand and pocket structures across different modeling setups. These curves reveal the proportion of samples below varying RMSD thresholds, offering a finer-grained perspective than summary statistics such as mean or median.
The first task evaluates the ability to predict ligand structures given both the apo and holo pocket conformations. UniGenX exhibits clear superiority across all metrics: the percentage of ligand predictions within 2Å RMSD improves from 1.27% to 21.66%, representing a 17-fold increase, while the median RMSD drops from 3.82Å to 2.37Å. Consistent gains are also observed at the 25th and 75th percentiles, as well as in mean RMSD, indicating not only improved accuracy but also greater robustness across diverse samples.
The second task, which requires predicting both the holo pocket and ligand coordinates from the apo structure alone, presents a more demanding setting that involves modeling pocket rearrangements—an essential aspect of induced fit. Here, the advantages of UniGenX are even more pronounced. For pocket prediction, UniGenX achieves a median RMSD of 1.75Å and a 2Å success rate of 69.43%, compared to 2.96Å and 12.10% for the baseline. For ligand prediction, the improvements are similarly substantial. When the pocket is provided (Pocket Gv), UniGenX reaches 24.84% within 2Å RMSD, outperforming the baseline’s 18.47%. In the more challenging case where the pocket must also be inferred (Pocket Not Gv), UniGenX achieves 15.92% within 2Å—an over 23-fold improvement compared to the baseline’s 0.64%.
Evaluation of the joint protein-ligand complex structure further underscores these gains: UniGenX reduces the mean RMSD from 4.20Å to 2.72Å. Taken together, these results demonstrate that UniGenX not only improves absolute prediction accuracy but also captures the coupled conformational dynamics of protein-ligand interactions. This ability to model joint structural distributions under dynamic settings is critical for advancing molecular design and docking beyond static approximations.
To further illustrate the qualitative fidelity of predicted structures, Figure 16 shows representative examples for two protein-ligand complexes (PDB IDs: 4KZQ and 2VVN) under both modeling conditions—with and without the holo pocket provided. In both cases, the predicted ligand structures (shown in cyan) exhibit strong agreement with the ground truth conformations, even when the pocket geometry must be inferred. These visual results reinforce the model’s ability to recover realistic and physically plausible docking configurations, aligning well with the quantitative improvements described above.
Figure 15: Cumulative Distribution of RMSD Values for Predicted Ligand and Pocket Structures.
Table 11: Evaluation of UniGenX (100M) versus 3DMolFormer(100M) on dynamic pocket-ligand prediction. RMSD percentiles, mean values, and the percentage of predictions within 2Å and 5Å are reported for pocket, ligand, and complex structures.
Percentiles $\downarrow$ % Below $\uparrow$
Method 25% 50% 75% Mean
3DMolFormer (100M)[33]
   Pocket + Ligand 3.03 3.82 4.66 4.20 1.27 82.17
   Pocket 2.34 2.96 4.06 3.54 12.10 85.99
   Ligand (Pocket Gv) 2.18 2.84 3.63 2.97 18.47 95.54
   Ligand (Pocket Not Gv) 4.15 5.22 6.60 5.41 0.64 43.31
UniGenX (100M)
   Pocket + Ligand 2.05 2.37 2.94 2.72 21.66 95.54
   Pocket 1.51 1.75 2.14 2.12 69.43 95.54
   Ligand (Pocket Gv) 2.01 2.57 3.29 2.73 24.84 98.09
   Ligand (Pocket Not Gv) 2.10 2.81 3.51 2.90 15.92 96.82
(a) 4KZQ Pocket given
(b) 2VVN Pocket given
(c) 4KZQ Pocket not given
(d) 2VVN Pocket not given
Figure 16: Examples of the docking predictions.
4.5 Unification Results of molecule and materials
@@
[…]
@@ 4.6 Intergrating with LLM Results
[…]
NatureLM[38] 61.78 0.044
UniGenX-NL(1B) $\mathbf{62.82}$ $\mathbf{0.042}$
Table 14: Results of UniGenX with capability of natural langauge on MP-20.
As shown in Table 14, our model UniGenX-NL(1B), which incorporates natural language understanding capabilities, not only preserves but also improves upon the performance of prior models such as NatureLM. Specifically, it achieves a higher matching rate (MR) of 63.10% and a lower RMSD of 0.040 Å, outperforming NatureLM on both metrics.
Table 15 show that comparing with UniGenX(100M), UniGenX-NL(1B) also performed better on QM9.
These results highlight the effectiveness of our approach in bridging linguistic understanding with structural prediction tasks. This also means that we can greatly enhance the model’s flexibility while maintaining its performance.
Dataset Large-scale QM9
[…]
Mean Median Mean Median Mean Median Mean Median
UniGenX(100M) 91.43 $\mathbf{100.00}$ 0.2186 $\mathbf{0.0818}$ 92.53 $\mathbf{100}$ 0.1441 0.0818
UniGenX–NL(1B) $\mathbf{92.30}$ $\mathbf{100.00}$ $\mathbf{0.1518}$ 0.0886 $\mathbf{92.88}$ $\mathbf{100}$ $\mathbf{0.1368}$ $\mathbf{0.0751}$
Table 15: Comparison between 100M model and Results of UniGenX with capability of natural langauge on QM9.
The results demonstrate that our model retains its linguistic capabilities without compromising accuracy. The synergistic effect of combining linguistic and numerical reasoning opens the door for developing unified models that can jointly process textual descriptions and scientific data. This paves the way for future applications such as text-guided molecule design, multimodal scientific assistants, and interactive generative systems in computational chemistry and materranh1
@@ 5 Conclusion
Appendix A Auto-regressive model, diffusion model and flow matching
A.1 Autoregressive Model
$\mathbf{x}={x_{1},x_{2},\ldots,x_{T}},$
the auto-regressive approach models the probability as
$P(\mathbf{x})=\prod_{t=1}^{T}P(x_{t}\mid x_{1},x_{2},\ldots,x_{t-1}),$
A.2 Diffusion Models
A.2.1 Diffusion Process
$\mathrm{d}x_{t}=f(x_{t},t)\mathrm{d}t+g(t)\mathrm{d}\bm{B}_{t},$ (7)
where $\bm{B}_{t}$ is a standard Wiener process, $f(\cdot,t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift coefficient of $x(t)$, and $g(\cdot):\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient of $x(t)$.
$p(x_{t})=\int_{x}p(x_{t}|x)p(x)dx,$ (9)
$x_{t}=\alpha_{t}x+\sigma_{t}\epsilon.$ (10)
Here $\alpha_{t}$ and $\sigma_{t}$ are linear and variance variable, depending on the time $t$; $\epsilon$ is a random variable. In the diffusion model, $\alpha_{t}$ and $\sigma_{t}$ are given by designed schemes for VE and VP cases [53]; the random variable $\epsilon$ is a standard Gaussian noise variable, i.e., $\epsilon\sim N(0,I)$. In our work, we use the Denoising Diffusion Probabilistic Model (DDPM) setting for training and the VP scheme for sampling [8, 52]. We also test the diffusion choice of Elucidated Diffusion Model (EDM) [54]. We obtained similar performance leveraging DDPM and EDM respectively, and listed the results as an ablation study in Appendix E.3.2.
A.2.2 Diffusion Targets
$p(x_{t}|x)=Ce^{-\frac{-\|x_{t}-\alpha_{t}x\|^{2}}{2\sigma_{t}^{2}}}.$ (11)
$s(x_{t})=\frac{\int_{x}\frac{\nabla_{x_{t}}p(x_{t}|x)}{p(x_{t}|x)}p(x_{t}|x)p(x)dx}{p(x_{t})}=\int_{x}\nabla_{x_{t}}\log p(x_{t}|x)p(x|x_{t})dx,$ (12)
where we finally obtain
${\theta}^{*}=\underset{{\theta}}{\arg\min}\,\mathbb{E}_{t\sim U[0,1]}\bigg{\{}\lambda(t)\,\mathbb{E}_{x}\,\mathbb{E}_{x_{t}\mid x}\big{[}\|{s}_{{\theta}}(x_{t},t)-\nabla_{x_{t}}\log p(x_{t}\mid x)\|_{2}^{2}\big{]}\bigg{\}},$
$s_{\theta^{*}}(x_{t})$ $=-\frac{\epsilon_{\theta^{*}}(x_{t})}{\sigma_{t}}.$
A.3 Connections of the Training Targets
@@
A.4 Diffusion model and flow matching
Theorem 1.
$\min_{Z_{\theta}\in\mathcal{V}}\|Z_{\theta}-X\|\,$,
where $\|X\|=\left(\int X^{2}\,dP\right)^{\frac{1}{2}}$.
$x_{t}=\alpha_{t}x+\sigma_{t}z,$ (19)
$\frac{dx_{t}}{dt}=\dot{\alpha}_{t}x+\dot{\sigma}_{t}z,$ (20)
@@ Appendix B Pseudocode
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
17:  $scores\leftarrow$ logits_warper(logits_processor($w_{gt},w_{logits[:-1]}),-1)$
@@
[…]
@@ C.1 Dictionary for Diverse Domains and Tasks
The vocabulary dictionary also incorporates special tokens, which are crucial for UniGenX’s flexibility as a next-token prediction model across diverse domains. These tokens serve as domain indicators, such as <<bof>> for material formulas and <<bos>> for molecule SMILES. Extending UniGenX to other domains, like proteins and DNA, requires adding corresponding tokens, such as <<boa>> and <<bod>>. This design empowers UniGenX to handle various domains within a unified sequence representation. Thus, these domain-specific special tokens are integral to the vocabulary dictionary.
Furthermore, special tokens facilitate different tasks. For structure generation, <<boc>> denotes coordinates. For conditional generation of sequences and structures, property tokens like <<bulk>> for bulk modulus and <<band>> for band gap are used. These tokens, followed by property values, precede the molecule sequence and coordinates. This allows UniGenX to generate both sequences and structures conditioned on specific properties. Therefore, the special tokens in the vocabulary dictionary provide UniGenX with flexibility across domains and tasks.
@@ C.5 Conditional Generation of Molecules
C.6 Conditional Generation of Materials
For conditional generation in the materials domain, we fine-tune a model that has been pretrained on large-scale materials data. The single-property datasets for bulk modulus and magnetization are taken directly from the [17]. For multi-property conditional generation, we curate a new dataset by filtering materials from the Materials Project and computing their properties using [19].
Each training sentence encodes a set of target properties through dedicated property tokens. Specifically, for a sample with $k$ conditioning attributes, the input sequence includes $k$ corresponding special tokens. The input sentence format is illustrated below:
<prop1> [prop1_val] <prop2> [prop2_val] <prop3> [prop3_val] <bos> [n * sites] <coord> [3 * lattice] [n * frac_coords] <eos>
C.7 Protein and Emulating MD equilibrium distributions
To train our foundational protein model, we constructed a sequence of amino acid chains (represented by their single-letter abbreviations) for each protein. For the rare unconventional amino acids in the ’Emulating MD equilibrium distributions’ section, we used the letter ’X’ as a substitute. After generation, a proprietary post-processing procedure was applied to ensure their consistency and validity.
The full input sequence is organized as follows:
[…]
For the evaluation metrics, we first partitioned the 2D TICA projections into a grid of $num\_bins\times num\_bins$ along the x and y axes. We then counted the number of points falling into each bin, which can be interpreted as an unnormalized probability distribution $P(x,y)$. The free energy was subsequently computed using the Boltzmann relation:
$F(x,y)=-k_{B}T\cdot ln(P(x,y)+\epsilon).$ (22)
To improve visualization, we offset the minimum free energy to zero and truncated extremely high energy values. This procedure yielded a more interpretable representation of the Free Energy Landscape.
C.8 Protein-Ligand Docking
For the protein-ligand docking task, we adopt a similar strategy to that used in the small molecule generation setting. Each small molecule is represented using its canonical SMILES string, while protein pockets are described at the full-atom level, with atom order aligned to the residue indexing in the original MD.hdf5 files provided by [34]. The corresponding coordinate tensors are arranged to follow this atom sequence exactly.
To enable a single model to handle both the ligand-only and joint pocket-ligand generation tasks, we design a unified input sequence structure that encodes all relevant components. The full input sequence is organized as follows:
<bos> <bopo> [n * pocket_atoms] <eopo> <boapc> [n * apo_coords] <eoapc> <bom> [cano_smiles] <eom> <bohpc> [n * holo_coords] <eohpc> <bomc> [m * lig_docked_coords] <eomc> <eos>
During inference, task-specific tokens are used to control which parts of the structure the model is expected to generate. For the first task—ligand generation conditioned on both apo and holo pockets—the input sequence is provided up to the <bomc> token, allowing the model to autoregressively infer the subsequent ligand coordinates. For the second task—joint prediction of the holo pocket and ligand given only the apo pocket and SMILES—the input sequence is truncated at <bohpc>, prompting the model to predict both the holo pocket and downstream ligand coordinates.
Since the primary learning objective is to recover 3D coordinates, training is supervised using only the diffusion loss, applied to the predicted atomic positions. Following the design of 3D-MolFormer, we apply coordinate scaling as part of preprocessing. All 3D coordinates are scaled down by a factor of 5 to facilitate more stable and effective learning.
C.9 EC Number-guided Protein Generation
For training data, we used all protein sequences in UniProt that have an assigned EC number and corresponding structural information available in the AlphaFold Database (AFDB), and with sequence lengths below 512 residues. This results in a dataset of approximately 20 million samples.
The first three digits of the EC number are decomposed into three separate numbers. For example, ”2.6.1.x” would be converted to 2 (ec number 1), 6 (ec number 2), and 1 (ec number 3).
<bos> <ec1> [ec number 1] <ec2> [ec number 2] <ec3> [ec number 3] [n * amino_acids] <coord> [n * coords] <eos>
C.10 Instruction Tuning with Natural Language Model